OpenStudy (anonymous):

Suppose that x and y vary inversely, and x = 3 when y = 10. Write the function that models the inverse variation.

OpenStudy (anonymous):

@sarahusher

OpenStudy (bahrom7893):

y = k(1/x) 10 = k(1/3)

OpenStudy (bahrom7893):

k = 3*10 = 30

OpenStudy (bahrom7893):

y = 30/x

OpenStudy (anonymous):

Can you help me with some more @bahrom7893 thanks

OpenStudy (bahrom7893):

Sorry, probably not, just follow along what I did.

OpenStudy (anonymous):

Its fine, but can you just try this one Suppose that y varies directly with x and inversely with z, and y = 50 when x = 15 and z = 3. Write the equation that models the relationship. Then find y when x = 8 and z = 4.

OpenStudy (anonymous):

@bahrom7893